LTO Report
Kitt Peak Nightly LTO Report For UT: 2025-04-03
Report Time (MST): 04:50
For access to all LTO-Reports go to: http://www-kpno.kpno.noirlab.edu/cgi-bin/ltoreport
4 Meter Weather Averages
From: UT = 01:00 (MST=6:00 PM) To: UT = 13:00 (MST=6:00 AM)
Elements | Mean | High | Low | AvDev |
Air Temperature (°C) | -0.5 | 1.6 | -1.3 | 0.6 |
Dewpoint Temperature (°C) | -7.4 | -3.6 | -12.9 | 2.1 |
Relative humidity (%) | 58 | 84 | 33 | 12.4 |
Wind Speed (mph) | 17.5 | 37.3 | 3.1 | 7.6 |
Wind Direction (°) | 214 | 4 | 184 | 0.8 |
4M Dome Temperature (°C) | 0.6 | 6.3 | -1.1 | 1.8 |
4M Dome Dewpoint (°C) | -5.9 | -3.4 | -9.4 | 1.5 |
4M Primary Mirror (°C) | 4.0 | 6.2 | 2.7 | 1.0 |
---|
Weather Comments: | A few light clouds at the beginning of the night gave way to mostly clear skies. Humidity ranged from 40 to 85%. Wind from the southwest until 0230 when it shifted to coming from the southeast. Wind speed fell from 35 to 5 mph. |
Recognized Hours: | ( 9 ) 18 degree twilight + 0.5 hr. |
Observer Information
Telescope | Observer | Prop. # | Instrument | Detector | Seeing |
4 Meter | LO: Elliott (on site); SO: Luo (remote, all night) | --- | DESI | N/A | 1.4 |
WIYN | Mintz, Ward (on site) | 2025A-130275 | HYDRA | STA1 | 1.5 |
2.1 Meter | No Observer | --- | N/A | N/A | N/A |
Feed | No Observer | --- | N/A | N/A | N/A |
0.9 Meter | No Observer | --- | N/A | N/A | N/A |
---|
Time Usage During The Night
Telescope | Hours Used | Night's Activity | | Hours Lost | Time Lost Due To: | Time of Night (MST) |
|
4Meter |
---|
3.29 | Scientific Research | | 0.59 | Weather | 19:00 - 20:00 |
| | | 5.00 | Dome Enclosure Problem | 00:00 - 05:30 |
| | | 0.12 | Misc. Problems | 22:00 - 22:30 |
|
WIYN |
---|
8.39 | Scientific Research | | 0.61 | Tele-Mechanical Problem | 20:30 - 21:30 |
|
2Meter |
---|
9.25 | Shutdown | | 0 | Weather | |
|
Feed |
---|
0 | Scientific Research | | 0 | Weather | |
|
W0_9m |
---|
9.25 | Shutdown | | 0 | Weather | |
Observing Status
4 Meter | DESI Survey Observing. Lost about 30 minutes at the beginning of the night due to high dust levels. Lost 7 minutes around 2200 to the VC caravan departing. Lost 5 hours at the end of the night due to a problem with dome slip rings. See trac ticket 1238 for details. Thanks to Bob Stupak for the phone assist! Seeing ranged from 1.2-1.6 arcsec as measured with DESI. Median seeing of 1.4 arcsec. |
WIYN | Hydra science observing. Lost 0.61 hours to MOS rotator issues. See WIYN Nightlog. No other time lost. Seeing ranged from 1 to 2 arcsec as estimated from Hydra FOPS. Median seeing of 1.5 arcsec. |
2.1 Meter | Not Scheduled |
Feed | Not Scheduled |
0.9 Meter | WIYN 0.9M shutdown until further notice. |
---|
Comments | A clear night with some technical challenges for the Mayall and WIYN. |
---|
Weather Plots




